# Gauge invariance and color charge fluctuations

### v1

@@ I Introduction
Understanding the nature of confinement has been a central theoretical problem in QCD from its earliest days[1]. Indeed, precisely what is meant by “confinement”[2] is not totally clear, beyond the very basic fact that QCD does not have isolated states with net color: there are no isolated quarks or gluons, so it is natural to think of the issue as one of color confinement. A related issue of phenomenological significance is that at sufficiently high temperatures, the thermodynamics of QCD acts like that of a plasma of quarks and gluons; this suggests that the quarks and gluons enter a deconfined quark-gluon plasma phase[3]. While Cabbibo and Parisi’s original idea that there the quark-gluon plasma is a distinct deconfined phase is not quite right, the notion of a quark-gluon plasma has become central in nuclear physics and has led to a massive experimental effort in heavy ion physics to determine its properties (for a review see [4]) . There have been numerous approaches to understanding the notion of confinement in QCD (or Yang-Mills theory) since the earliest days of the theory [^1] These include approaches aimed at understanding issues of confinement via a linear rising potential and the area law of the Wilson loop[1, 5], the role of center symmetry and the role of center symmetry breaking and the Polyakov loop[6, 7], approaches based on ’t Hooft loops, the vertex free energy, or center vertices [8, 9], dual superconductors [10, 11] or the Kugo-Ojima approach[12]. These and other approaches were extensively studied in the decades since their formulation; the state of the art circa 2010 was reviewed in ref. [2]. Theoretical attempts to understand confinement are ongoing and new approaches are still being developed—such as a recent scheme describing deconfinement via percolation of electric center fluxes[13]. But it is probably fair to say that there is no real consensus on how confinement works or even what it means to understand confinement.
@@
However, as will be shown in this paper, despite the non-gauge-invariant nature of the color charge and the triviality of the color charge matrix elements, there are gauge invariant —- and hence physical—measures of color charge fluctuations in spatial regions. Even though the net global color charge of any physical state is zero, this does not rule out, nonzero color charges in different spatial regions with correlations yielding a vanishing net color but nonzero color fluctuations. An analogy with isopin might be helpful. A theory with exact isospin can have a state $|\psi\rangle$ that is isosinglet ($I^{2}=0$) but which has two spatial regions, each with isospin content. If the isospin of the two regions are $\vec{I}_{1}$ and $\vec{I}_{2}$ with $\vec{I}=\vec{I}_{1}+\vec{I}_{2}$, then the Wigner-Eckert theorem implies that $\langle\psi|\vec{I}_{1}|\psi\rangle=\langle\psi|\vec{I}_{2}|\psi\rangle=0$. However, this does not imply that there is no isospin in each region, but rather that the isospin in each region can fluctatuate in a correlated way with $\langle\psi|\vec{I}_{1}^{2}|\psi\rangle=\langle\psi|\vec{I}_{2}^{2}|\psi\rangle=-\langle\psi|\vec{I}_{1}\cdot\vec{I}_{2}+\vec{I}_{2}\cdot\vec{I}_{1}|\psi\rangle$. That is, while the expectation value of every component of the isospin in each regions is zero, the fluctuations of isospin need not be.
@@
is a color singlet (and, thus not forced to vanish by the Wigner-Eckert theorem); its expectation value yields the color fluctuations. However, unlike with isospin there is also local gauge symmetry and ${\cal Q}^{2}_{R}$ is not gauge invariant. It is often assumed that unless an operator is gauge invariant, it is unphysical, which might suggest that ${\cal Q}^{2}_{R}$ is not physically relevant. However, the assumption that non-gauge-invariant operators are necessarily unphysical is not quite right.
@@
$\begin{split}&\langle{\cal Q}^{2}_{R}\rangle_{\rho_{\rm phys}}=\\ &\left(\frac{a_{l}}{g(a_{l})}\right)^{2}\,{\rm Tr}\left[\rho_{\rm phys}\int_{\partial R}d^{2}x\left(\hat{n}(\vec{x})\cdot\vec{E}^{a}(\vec{x})\right)^{2}\right]\;,\end{split}$ (5b)
@@
$\begin{split}&\langle{\cal Q}^{2}_{R}\rangle_{\rho_{\rm phys}}=\\ &\left(\frac{a_{l}}{g(a_{l})}\right)^{2}\sum_{k\in\partial R}\sum_{f}\sum_{a}{\rm Tr}\left[\rho_{\rm phys}\left(\vec{E}_{k}^{a}\cdot\hat{n}_{k}^{f}\right)^{2}\right]\end{split}$ (6)
where $k$ represents lattice sites on the boundary of the region and $f$ a face on that cell which forms the boundary (so that $\hat{n}_{f}^{k}$ represents the unit vector orthogonal to that face). The analysis goes through, mutatis mutandis, for other choices of regulator. The righthand side of Eq.  (6) is gauge invariant; thus, $\langle{\cal Q}^{2}_{R}\rangle_{\rho_{\rm phys}}$ is gauge-invariant even though ${\cal Q}^{2}_{R}$ is not. This is a principal result of this paper.
@@
Figure 1: Each block is cubic and with the same volume. Region A is a single block, Region B has two disconnected blocks, Region C has two connected blocks with one common face, region D has 7 connected blocks with a total of 7 faces shared by two distinct blocks .
@@
$\begin{split}&V_{A}:V_{B}:V_{c}:V_{D}=1:2:2:7\;\;{\rm while}\\ &\langle Q^{2}_{R_{A}}\rangle_{\rm vac}:\langle Q^{2}_{R_{B}}\rangle_{\rm vac}:\langle Q^{2}_{R_{C}}\rangle_{\rm vac}:\langle Q^{2}_{R_{D}}\rangle_{\rm vac}\\ &=1:2:\frac{5}{3}:\frac{14}{3}\;.\end{split}$ (9)
As noted earlier, the fluctuations do not scale with volume: region D is 7 times the volume of region A, but the color charge fluctuations are only $\frac{14}{3}$ times larger. Regions B and C illustrate a related peculiarity: the fluctuations in region B are larger than in C by a factor of $\frac{6}{5}$ due to the gap in the disconnected regions of $B$. That the ratio remains $\frac{6}{5}$ even if the gap size is altered but if the regions merge, the ratio becomes unity.
@@
As a final example, consider the ground state of Yang-Mills theory with two well-separated fundamental color charges (ie. arbitrarily heavy quarks) added to the system in a gauge-invariant way via a Wilson line. The behavior of such a system has long been known[1, 5, 19, 20, 2]: close to the sources, the color electric fields are Coulomb-like—$\left(\vec{E}^{2}-\langle\vec{E}^{2}\rangle_{\rm vac}\right)$ falls off as $1/r^{4}$—up logarithmic corrections; well away from the sources, a flux tube (responsible for a linear rising potential) forms. A small spherical region centered on one of the charges has $F_{R;\rho_{\rm phys}}$ falling off as $1/r^{2}$ (up to logarithmic corrections). This might appear to suggest that $F_{R;\rho_{\rm phys}}$ represents a color charge localized at the source with some form of color screening. However, the interpretation of $F_{R;\rho_{\rm phys}}$ is not viable. Considers cylindrical region coaxial with the flux tube. Recall that well away from well-separated charges, the flux tube has a more or less constant width with a color electric field strength that is approximately independent of the distance to the charges. For a cylindrical region coaxial with the flux tube and wide enough to fully contain it, one sees that $F_{R;\rho_{\rm phys}}$ is non-zero and has a value that is approximately independent of the length of the cylinder and where it is placed between the sources. This implies shifts in the color fluctuations from the vacuum in regions well away from the sources is nonzero and independent of the size of the region; this looks nothing like color charges localized at the source with some form of color screening.

### v3

@@ I Introduction
Understanding the nature of confinement has been a central theoretical problem in QCD from its earliest days[1]. Indeed, precisely what is meant by “confinement”[2] is not totally clear, beyond the very basic fact that QCD does not have isolated states with net color: there are no isolated quarks or gluons, so it is natural to think of the issue as one of color confinement. A related issue of phenomenological significance is that at sufficiently high temperatures, the thermodynamics of QCD acts like that of a plasma of quarks and gluons; this suggests that the quarks and gluons enter a deconfined quark-gluon plasma phase[3]. While Cabbibo and Parisi’s original idea that there the quark-gluon plasma is a distinct deconfined phase is not quite right, the notion of a quark-gluon plasma has become central in nuclear physics and has led to a massive experimental effort in heavy ion physics to determine its properties (for a review see [4]) . There have been numerous approaches to understanding the notion of confinement in QCD (or Yang-Mills theory) since the earliest days of the theory [^1] These include approaches aimed at understanding issues of confinement via a linear rising potential and the area law of the Wilson loop[1, 5], the role of center symmetry and the role of center symmetry breaking and the Polyakov loop[6, 7], approaches based on ’t Hooft loops, the vertex free energy, or center vertices [8, 9], dual superconductors [10, 11] or the Kugo-Ojima approach[12]. These and other approaches were extensively studied in the decades since their formulation; the state of the art circa 2010 was reviewed in ref. [2]. Theoretical attempts to understand confinement are ongoing and new approaches are still being developed—such as a recent scheme describing deconfinement via percolation of electric center fluxes[13].
It is probably fair to say that there is no real consensus on how confinement works or even what it means to understand confinement. However, the most basic notion of confinement is some variant of the idea that particles carrying a color charge cannot be isolated. It is quite easy to show that gauge invariance requires the expectation value of every component of the net color charge in any spatial region to be zero for any physical state in the theory. However, confinement in the most basic sense clearly means something more than that: imagine a hypothetical isolated quark or gluon in a state where the direction in color-space has quantum fluctuations such that average color in any direction is zero but where the square of the of the net color charge—a measure of the fluctuations—in a spatial region surrounding the quark to be nonzero. Such a state should not be possible in the physical spectrum of a confined theory. Thus, understanding color charge fluctuations may be a key to gaining insights on confinement [^2] Understanding color charge fluctuations is of importance for other aspects of strong interaction physics. See for example ref. [14].
@@
However, as will be shown in this paper, despite the non-gauge-invariant nature of the color charge and the triviality of the color charge matrix elements, there are gauge invariant—-and hence physical—measures of color charge fluctuations in spatial regions. Even though the net global color charge of any physical state is zero, this does not rule out, nonzero color charges in different spatial regions with correlations yielding a vanishing net color but nonzero color fluctuations. An analogy with isopin might be helpful. A theory with exact isospin can have a state $|\psi\rangle$ that is isosinglet ($I^{2}=0$) but which has two spatial regions, each with isospin content. If the isospin of the two regions are $\vec{I}_{1}$ and $\vec{I}_{2}$ with $\vec{I}=\vec{I}_{1}+\vec{I}_{2}$, then the Wigner-Eckert theorem implies that $\langle\psi|\vec{I}_{1}|\psi\rangle=\langle\psi|\vec{I}_{2}|\psi\rangle=0$. However, this does not imply that there is no isospin in each region, but rather that the isospin in each region can fluctuate in a correlated way with $\langle\psi|\vec{I}_{1}^{2}|\psi\rangle=\langle\psi|\vec{I}_{2}^{2}|\psi\rangle=-\langle\psi|\vec{I}_{1}\cdot\vec{I}_{2}+\vec{I}_{2}\cdot\vec{I}_{1}|\psi\rangle$. That is, while the expectation value of every component of the isospin in each regions is zero, the fluctuations of isospin need not be.
@@
is a color singlet (and, thus not forced to vanish by the Wigner-Eckert theorem); its expectation value yields the color fluctuations. However, unlike with isospin there is also local gauge symmetry and ${\cal Q}^{2}_{R}$ is not gauge invariant [^3] An easy way to see that it is not gauge invariant is to consider a configuration in which the color charge density is non-zero only in two spatially disjoint regions denoted $1$ and $2$; in both regions the color charge density points in a single color direction $b$: $\rho^{a}_{\rm color}=\rho(\vec{x})\delta_{ab}$ with $\rho(\vec{x})=0$ unless $\vec{x}$ is in region 1 or 2. Consider a region $R$ that fully contains both regions 1 and 2; then ${\cal Q}^{2}_{R}$ has contributions from region 1, region 2 and a non-zero cross term. Next consider a gauge transformation which is unity over region 1 and is a constant color rotation over region 2 to a direction orthogonal to $b$. The contribution to ${\cal Q}^{2}_{R}$ from region 1 and 2 will remain unchanged under this transformation, but the cross term vanishes.. It is often assumed that unless an operator is gauge invariant, it is unphysical, which might suggest that ${\cal Q}^{2}_{R}$ is not physically relevant. However, the assumption that non-gauge-invariant operators are necessarily unphysical is not quite right.
@@
$\begin{split}&\langle{\cal Q}^{2}_{R}\rangle_{\rho_{\rm phys}}=\\ &\left(\frac{a_{l}}{g(a_{l})}\right)^{2}\,{\rm Tr}\left[\rho_{\rm phys}\int_{\partial R}\!\!\!d^{2}x\sum_{a}\left(\hat{n}(\vec{x})\cdot\vec{E}^{a}(\vec{x})\right)^{2}\right]\ ,\end{split}$ (5b)
@@
$\begin{split}&\langle{\cal Q}^{2}_{R}\rangle_{\rho_{\rm phys}}=\\ &\left(\frac{a_{l}}{g(a_{l})}\right)^{2}{\rm Tr}\left[\rho_{\rm phys}\sum_{k\in\partial R}\sum_{f}\sum_{a}\left(\vec{E}_{k}^{a}\cdot\hat{n}_{k}^{f}\right)^{2}\right]\end{split}$ (6)
where $k$ represents lattice sites on the boundary of the region and $f$ a face on that cell which forms the boundary (so that $\hat{n}_{f}^{k}$ represents the unit vector orthogonal to that face). The analysis goes through, mutatis mutandis, for other choices of regulator. The operator $\sum_{k\in\partial R}\sum_{f}\sum_{a}\left(\vec{E}_{k}^{a}\cdot\hat{n}_{k}^{f}\right)^{2}$ on the right-hand side of Eq.  (6) is gauge invariant; thus, the value of $\langle{\cal Q}^{2}_{R}\rangle_{\rho_{\rm phys}}$ is unchanged by gauge transformations on the operator—it is gauge invariant—even though the operator on which it is based, ${\cal Q}^{2}_{R}$, is not. This is a principal result of this paper.
@@
Figure 1: Each block is cubic and with the same volume. Region A is a single block, Region B has two disconnected blocks, Region C has two connected blocks with one common face, region D has 7 connected blocks with a total of 8 faces shared by two distinct blocks .
@@
$\begin{split}&V_{A}:V_{B}:V_{c}:V_{D}=1:2:2:7\;\;{\rm while}\\ &\langle Q^{2}_{R_{A}}\rangle_{\rm vac}:\langle Q^{2}_{R_{B}}\rangle_{\rm vac}:\langle Q^{2}_{R_{C}}\rangle_{\rm vac}:\langle Q^{2}_{R_{D}}\rangle_{\rm vac}\\ &=1:2:\frac{5}{3}:\frac{13}{3}\;.\end{split}$ (9)
As noted earlier, the fluctuations do not scale with volume: region D is 7 times the volume of region A, but the color charge fluctuations are only $\frac{13}{3}$ times larger. Regions B and C illustrate a related peculiarity: the fluctuations in region B are larger than in C by a factor of $\frac{6}{5}$ due to the gap in the disconnected regions of $B$. That the ratio remains $\frac{6}{5}$ even if the gap size is altered but if the regions merge, the ratio becomes unity.
@@
[…]